OpenStudy (svnnynight9):

Trigonometry sec2x=? check all that apply A. 1/cos^2x-sin^2x B.1/1-2sin^2x C.1/sin^2x-cos^2x D.1/2sin^2x-1

OpenStudy (math&ing001):

cos(2x)=cos^2(x)-sin^2(x)=1-2sin^2(x)=2cos^2(x)-1 So \[sec(2x)=\frac{1}{cos^2(x)-sin^2(x)}=\frac{1}{1-2sin^2(x)}=\frac{1}{2cos^2(x)-1}\]

OpenStudy (svnnynight9):

Thanks

OpenStudy (math&ing001):

NP